# A Comprehensive Survey on GNN Characterization

### v2

@@ Abstract
Characterizing graph neural networks (GNNs) is essential for identifying performance bottlenecks and facilitating their deployment. Despite substantial work in this area, a comprehensive survey on GNN characterization is lacking. This work presents a comprehensive survey, proposing a triple-level classification method to categorize, summarize, and compare existing efforts. In addition, we identify promising future directions for GNN characterization. Our survey aims to help scholars systematically understand GNN performance bottlenecks and patterns from a computer architecture perspective, contributing to more efficient GNN execution.
@@
Graph neural network, Characterization, Quantitative analysis, Compute pattern, Memory pattern.
@@ 1 Introduction
Graph neural networks (GNNs) are a type of neural network specifically designed to process graph-structured data, where entities are represented as vertices and their relationships as edges. Unlike traditional deep neural networks (DNNs) that typically operate on grid-structured data such as images, GNNs leverage the connection patterns within graphs to perform tasks such as vertex classification, link prediction, and graph classification [51, 29, 38]. By iteratively aggregating and transforming feature information from neighboring vertices, GNNs capture the information of both the local neighborhood and the global graph structure. This ability to naturally handle the irregular structure of graphs gives GNNs a significant advantage over traditional DNNs, offering flexibility and adaptability in tasks involving relational data [7]. Consequently, GNNs have become powerful tools for applications including knowledge reasoning, recommendation systems, and traffic forecasting [51, 29, 38].
@@
Characterizing and understanding GNNs is crucial for identifying their performance bottlenecks and optimizing their deployment across various applications. This process involves systematically measuring and evaluating different aspects of GNNs from a computer architecture perspective, such as compute pattern, memory pattern, and communication pattern [55, 58, 69]. These insights provide essential foundational knowledge and theoretical guidance for enhancing the performance and efficiency of GNN execution. Consequently, researchers and engineers can propose effective software and hardware optimizations, such as strategies for eliminating redundant computations and exploiting parallelism, to address performance bottlenecks and facilitate the widespread adoption of GNNs. Substantial work has already been done in this area [55, 58, 69, 4], highlighting the importance of such characterization in improving GNN implementations and hardware performance.
@@
However, there is currently a lack of comprehensive surveys specifically focused on GNN characterization. Recent surveys related to GNNs primarily concentrate on algorithmic features [51, 24, 50, 34, 32, 6], algorithmic and hardware acceleration techniques [33, 1, 27], and distributed training [29, 38]. A consolidated survey systematically reviewing the characterization efforts for GNNs is notably absent. Such a survey is essential for systematically organizing, summarizing, and comparing existing efforts and findings in GNN characterization, providing readers with a comprehensive understanding and overview. This can prevent redundant work and consolidate a vast amount of findings in one place, enabling researchers and engineers to quickly access relevant knowledge. For newcomers to the GNN research field, such a survey can serve as a comprehensive learning resource, helping them quickly grasp key findings and recent advancements.
@@
In this work, we present a comprehensive survey on GNN characterization from the perspective of computer architecture. Initially, we propose a triple-level classification method that coordinates GNN model types with computer architecture perspectives. This approach first categorizes previous efforts according to the types of GNNs and then systematically summarizes the characterization findings based on their computer architecture perspectives and further subcategories. Additionally, we compare the findings across different types of GNNs to provide a holistic understanding from the perspective of computer architecture. This comparison sheds light on the different execution behaviors and characteristics of various GNN types, offering valuable insights into optimizing GNN implementations. Finally, we identify promising directions for future research in GNN characterization.
@@
In summary, by consolidating these characterization findings, our survey aims to help scholars quickly, systematically, and comprehensively understand the performance bottlenecks and various program patterns of GNNs from the viewpoint of computer architecture. This, in turn, contributes to the development of more efficient GNN implementations across diverse applications and hardware platforms and guides hardware designs for GNNs.
The organization of this survey is shown in Table I.
@@
Taxonomy Principle of Taxonomy § 3.1
@@ 2.1 Graphs and GNNs
Static heterogeneous GNNs (SHeGNNs) are specifically designed to handle static heterogeneous graphs, as shown in Fig. 1(c). They typically involve four primary execution stages: semantic graph build, feature projection, neighbor aggregation, and semantic fusion [60]. The semantic graph build stage partitions the heterogeneous graph into multiple semantic graphs based on a set of metapaths, with each semantic graph corresponding to a specific metapath. The feature projection and neighbor aggregation stages perform SHoGNN operations independently within each semantic graph, projecting vertex features and aggregating neighbor information. Finally, the semantic fusion stage fuses the aggregated results from all semantic graphs, integrating diverse types of high-order relationships to form comprehensive vertex embeddings.
@@ 2.3 GNN Characterization
Figure 3: Principle of taxonomy.
@@ 3 Taxonomy
3.1 Principle of Taxonomy
The principle of our taxonomy, as shown in Fig. 3, first classifies previous efforts based on their focus on model type. It then categorizes their findings according to dual-level architectural perspectives. This taxonomy is proposed because previous efforts often characterize GNNs from several main perspectives: 1) time (T), to identify hotspots of time and operation, and 2) compute (C), memory (M), communication (CO), parallelism (P), and execution (E), to deeply identify performance bottlenecks and discover opportunities for optimizing performance and efficiency.
Time. Analyzing the time percentage spent in different stages and operations reveals performance bottlenecks, making it the first step in optimization. Understanding the relationship between time and graph characteristics enables effective preprocessing and model implementation, while the relationship between time and model setup helps balance accuracy and speed. Additionally, understanding time complexity provides a theoretical basis for addressing bottlenecks.
@@ 3.2 Overview of Categorization
Table II presents a categorization overview of the latest efforts on GNN characterization using our proposed taxonomy. In addition, we include supplementary information in the table to offer a comprehensive review of these studies. The column labeled phase includes training and inference phases. The column labeled software framework lists the software frameworks used for characterization in the corresponding work. Two widely used software frameworks are PyG [12] and DGL [44]. The column of hardware platform enumerates the hardware platforms utilized to run the software frameworks in the corresponding work. “+” denotes that the hardware platforms include two or more components, such as CPU + GPU. “—” indicates that these efforts do not provide these details.
@@
[31] SHoGNN T, CO, P Training DGL, Euler [3] V100 GPUs
@@
[66] DHoGNN C, M, P Inference Titan Xp GPU
[10] DHoGNN C, M, P Inference
@@
[42] DHoGNN M, CO, P Training PyGT [37] CPU + V100 GPU
@@
[58] SHeGNN T, C, M, P, E Inference DGL T4 GPU
[53] SHeGNN T, M, P, E Inference PyG T4 GPU
@@ 4.1 Findings on Time Perspective
Time Percentage. T-1 During SHoGNN training, the first layer consumes significantly more time than the second layer in model computation due to dimensionality reduction of vertex features and aggregation from more neighbors in the first layer [46]. T-2 During SHoGNN inference, both the gather and apply_vertex substages can dominate the execution time of each layer, varying across different datasets and models due to factors such as the variable dimensions of input and output vertex features [56]. T-3 Similarly, during SHoGNN training, the aggregation stage typically dominates the execution time of each layer in model computation, except in models with high computational complexity in the apply_vertex substage or graphs with low average vertex degrees [46]. T-4 Edge-related computations occupy most of the execution time of SHoGNN training and inference in most models [46]. For models with high edge computational complexity, most of time is spent on the messaging function for each edge. Otherwise, the aggregation of messages from all edges consume most of the time. T-5 There is no fixed hotspot in SHoGNN training, as substage time distribution varies greatly among models [64, 4].
T-6 Sampling in mini-batch training and inference of SHoGNNs constitutes a significant portion of total time [22, 26, 46, 19, 31, 2]. Sampling, typically performed on the CPU, can take up to 62% of an epoch’s time [22]. This bottleneck worsens if the CPU is attached to multiple GPUs and cannot produce enough samples to saturate them.
T-7 Mini-batch training for SHoGNNs faces time-consuming feature collection and data loading from CPU to GPU, particularly for loading vertex features [30, 28, 2]. The study [30] observed that 74% of time is spent on feature collection and data loading. This is mainly due to the large size of graph data, resulting in longer transfer time than model computation time. Moreover, using multiple GPUs within a single physical machine to speed up training proportionally increases the demand for data samples loaded from CPU to GPU. T-8 Furthermore, feature collection is CPU-intensive and takes significantly longer than CPU-GPU data movement [30]. When multiple GPUs are used, concurrent workers collecting features compete with samplers for CPU resources, resulting in an 88% increase in sampling time and a 59% increase in feature collection time compared to the single-GPU scenario [30]. T-9 Data loading time for training graph classification models constitutes a significant proportion of total training time [48]. Unlike the general data loading focus in T-4.1, T-4.1 specifically focuses on batching a set of small graphs into a single large batch.
T-10 The gradient synchronization in model computation becomes more time-consuming as the number of GPUs increases [28]. Unlike DNNs, GNNs have smaller model sizes with fewer layers and shared weights across all vertices. While the gradient synchronization is expected to occupy a small part of the execution time, it takes up a larger proportion as GPU count increases due to CPU-side imbalances, causing staggered GPU computations.
Relation between Time and Graph Characteristics. T-11 During SHoGNN inference, the execution time of the gather and apply_vertex substages in each layer increases linearly with the dimension of the output hidden features [55]. T-12 Additionally, the execution time of the apply_vertex substage is almost proportional to the dimension of the input vertex features [55]. T-13 Similarly to T-4.1, which focuses on SHoGNN inference, the execution time of the scatter, apply_edge, gather, and apply_vertex substages during SHoGNN training increases linearly with the dimension of the output hidden features [46]. T-14 The training and inference time of a layer is primarily affected by the dimensions of the input and output hidden features, increasing linearly with these dimensions [46]. This is consistent with T-4.1, T-4.1, and T-4.1. T-15 The time of the aggregation stage in a layer during SHoGNN training increases linearly with the average vertex degree of the graph [46]. T-16 The average vertex degree of the graph affects the proportion of execution time for scatter, apply_edge, and gather substages in the total training time [46].
Relation between Time and Model Setup. T-17 Fixing other hyperparameters, each individual hyperparameter affects the training and inference time of a layer linearly [46]. T-18 Sampling reduces training time per batch only when the batch size is very small. As batch size increases, the overhead of sampling and data loading causes the training time to exceed that of full-batch training [46].
@@
Time Complexity. T-19 The training time is related to the model’s time complexity. T-20 The time complexity of the message function used in the apply_edge substage affects the execution time of the aggregation stage [46, 48]. T-21 The time complexity for each vertex is affected by the dimensions of the input and output hidden features ($d_{in}$ and $d_{out}$) and the dimensions of the model parameters, such as the number of heads ($K$) in GAT [41] [46]. This is consistent with T-4.1, T-4.1, T-4.1, and T-4.1. Accordingly, typical SHoGNNs are classified into four quadrants based on their time complexity of aggregation and combination, as shown in Fig. 4 [46]. T-22 All stochastic gradient descent based training algorithms suffer from exponential time complexity with respect to the number of layers [11, 48].
@@ 4.2 Findings on Compute Perspective
C-5 During SHoGNN training, computation is dominated by integer operations [4]. Results show that, on average, 64% of the executed instructions are integer instructions, while only 28.7% are floating-point instructions [4]. C-6 The scatter substage primarily involves data movement and address calculations for data accesses, preparing data for the subsequent substage [64, 40]. This explains the high proportion of integer operations observed in C-4.2.

### v3

# Survey on Characterizing and Understanding GNNs from a Computer Architecture Perspective

@@ Abstract
Characterizing and understanding graph neural networks (GNNs) is essential for identifying performance bottlenecks and facilitating their deployment in parallel and distributed systems. Despite substantial work in this area, a comprehensive survey on characterizing and understanding GNNs from a computer architecture perspective is lacking. This work presents a comprehensive survey, proposing a triple-level classification method to categorize, summarize, and compare existing efforts, particularly focusing on their implications for parallel architectures and distributed systems. We identify promising future directions for GNN characterization that align with the challenges of optimizing hardware and software in parallel and distributed systems. Our survey aims to help scholars systematically understand GNN performance bottlenecks and execution patterns from a computer architecture perspective, thereby contributing to the development of more efficient GNN implementations across diverse parallel architectures and distributed systems.
@@
Graph neural network, Characterization, Execution patterns, Performance bottlenecks, Parallel Architectures.
@@ 1 Introduction
Graph neural networks (GNNs) are a type of neural network specifically designed to process graph-structured data, where entities are represented as vertices and their relationships as edges. Unlike traditional deep neural networks (DNNs) that typically operate on grid-structured data such as images, GNNs leverage the connection patterns within graphs to perform tasks such as vertex classification, link prediction, and graph classification [50, 30, 38]. By iteratively aggregating and transforming feature information from neighboring vertices, GNNs capture the information of both the local neighborhood and the global graph structure. This ability to naturally handle the irregular structure of graphs gives GNNs a significant advantage over traditional DNNs, offering flexibility and adaptability in tasks involving relational data [7]. Consequently, GNNs have become powerful tools for applications including knowledge reasoning, recommendation systems, and traffic forecasting [50, 30, 38, 25, 26].
@@
Characterizing and understanding GNNs is crucial for identifying their performance bottlenecks and optimizing their deployment across various applications, particularly within parallel architectures and distributed systems. This process involves systematically measuring and evaluating different aspects of GNNs from a computer architecture perspective, such as compute pattern, memory pattern, and communication pattern [53, 56, 66]. These insights provide essential foundational knowledge and theoretical guidance for enhancing the performance and efficiency of GNN execution on parallel architectures and distributed systems. Consequently, researchers and engineers can propose effective software and hardware optimizations, such as strategies for eliminating redundant computations and exploiting parallelism, to address performance bottlenecks and facilitate the widespread adoption of GNNs. Substantial work has already been done in this area [53, 56, 66, 4], highlighting the importance of such characterization in improving GNN implementations and performance within the context of parallel architectures and distributed systems.
@@
However, there is currently a lack of comprehensive surveys specifically focused on GNN characterization. Recent surveys related to GNNs primarily concentrate on algorithmic features [50, 23, 49, 35, 33, 6], algorithmic and hardware acceleration techniques [34, 1, 28], and distributed training [30, 38]. A consolidated survey systematically reviewing the characterization efforts for GNNs is notably absent. Such a survey is essential for systematically organizing, summarizing, and comparing existing efforts and findings on characterizing and understanding GNNs, providing readers with a comprehensive understanding of their implications for parallel and distributed architectures. This can prevent redundant work and consolidate a vast amount of findings in one place, enabling researchers and engineers to quickly access relevant knowledge. For newcomers to the GNN research field, such a survey can serve as a comprehensive learning resource, helping them quickly grasp key findings and recent advancements.
@@
In this work, we present a comprehensive survey on characterizing and understanding GNNs from the perspective of computer architecture, with a particular focus on their implications for parallel architectures and distributed systems. Initially, we propose a triple-level classification method that coordinates GNN model types with computer architecture perspectives. This approach first categorizes previous efforts according to the types of GNNs and then systematically summarizes the characterization findings based on their computer architecture perspectives and further subcategories. Additionally, we compare the findings across different types of GNNs to provide a holistic understanding from the perspective of computer architecture. This comparison sheds light on the different execution behaviors and characteristics of various GNN types, offering valuable insights into optimizing GNN implementations for parallel architectures and distributed systems. Finally, we identify promising directions for future research in GNN characterization, particularly in the context of parallel architectures and distributed systems.
Our contributions are summarized as follows:
  • $\bullet$
    A triple-level classification framework for categorizing and analyzing efforts to characterize and understand GNNs from a computer architecture perspective.
  • $\bullet$
    A comprehensive comparison of findings across different types of GNNs, providing a holistic understanding through the lens of computer architecture.
  • $\bullet$
    A discussion of promising future research directions in GNN characterization.
@@
In summary, by consolidating these characterization findings, our survey aims to assist scholars in quickly, systematically, and comprehensively understanding the performance bottlenecks and diverse execution patterns of GNNs from the perspective of computer architecture. This understanding, in turn, contributes to the development of more efficient GNN implementations across a wide range of applications and parallel architectures, while also guiding hardware design for GNNs in parallel and distributed systems. The organization of this survey is shown in Table I.
@@
Taxonomy Methodology of Taxonomy § 3.1
@@ 2.1 Graphs and GNNs
Static heterogeneous GNNs (SHeGNNs) are specifically designed to handle static heterogeneous graphs, as shown in Fig. 1(c). They typically involve four primary execution stages: semantic graph build, feature projection, neighbor aggregation, and semantic fusion. The semantic graph build stage partitions the heterogeneous graph into multiple semantic graphs based on a set of metapaths, with each semantic graph corresponding to a specific metapath. The feature projection and neighbor aggregation stages perform SHoGNN operations independently within each semantic graph, projecting vertex features and aggregating neighbor information. Finally, the semantic fusion stage fuses the aggregated results from all semantic graphs, integrating diverse types of high-order relationships to form comprehensive vertex embeddings.
@@
The application of different GNN models is guided by their respective features, as described above. SHoGNNs are particularly suitable for applications where the graph structure does not change over time, such as social network analysis or recommendation systems that operate on a fixed set of relationships. DHoGNNs, on the other hand, are more appropriate for scenarios involving temporal changes in graph data, such as traffic prediction or stock market forecasting, where both graph structure and vertex features evolve over time. SHeGNNs can be effectively applied in domains where graphs contain diverse vertex and edge types, such as knowledge graphs or multi-modal data integration tasks. Meanwhile, DHeGNNs are suited for applications like real-time event tracking or evolving multi-relational data, where both the structure and the type of relationships between vertices vary dynamically over time.
@@ 2.3 GNN Characterization
Figure 3: Methodology of taxonomy.
@@ 3 Taxonomy
3.1 Methodology of Taxonomy
The methodology for classifying and analyzing prior efforts is built upon a systematic taxonomy, as shown in Fig. 3. It uses GNN model types and architecture perspectives as criteria for categorization and involves two key steps:
  • $\bullet$
    Step 1: Classification by GNN model type, including static homogeneous, dynamic homogeneous, static heterogeneous, and dynamic heterogeneous GNNs. This step ensures a clear distinction of efforts and findings across different GNN model types.
  • $\bullet$
    Step 2: Categorization by six architecture perspectives—Time (T), Parallelism (P), Compute (C), Memory (M), Communication (CO), and Execution (E). This step enables a deeper identification of performance bottlenecks, uncovers optimization opportunities, and distinctly highlights the focus of various findings.
This methodology builds on prior studies that typically focus on a specific GNN model type and emphasize the significance of the following architecture perspectives in diagnosing performance challenges in parallel and distributed systems.
Time. Understanding time complexity provides a theoretical basis for addressing bottlenecks. Analyzing the time percentage spent in different stages and operations reveals performance bottlenecks, making it the first step in optimization. Understanding the relationship between time and graph characteristics enables effective preprocessing and model implementation, while the relationship between time and model setup helps balance accuracy and speed.
@@ 3.2 Overview of Categorization
Table II presents a categorization overview of the latest efforts on GNN characterization using our proposed taxonomy. In addition, we include supplementary information in the table to offer a comprehensive review of these studies. The column labeled phase includes training and inference phases. The column labeled software framework lists the software frameworks used for characterization in the corresponding work. Two widely used software frameworks are PyG [12] and DGL [44]. The column of hardware platform enumerates the hardware platforms utilized to run the software frameworks in the corresponding work. “+” denotes that the hardware platforms include two or more components, such as CPU + GPU. “—” indicates that these efforts do not provide these details. Guided by our taxonomy methodology, this table enables a comparative analysis of prior efforts across different GNN model types, architecture perspectives, phases, software frameworks, and hardware platforms.
@@
[32] SHoGNN P, T, CO Training DGL, Euler [3] V100 GPUs
@@
[63] DHoGNN P, C, M Inference Titan Xp GPU
[10] DHoGNN P, C, M Inference
@@
[42] DHoGNN P, M, CO Training PyGT [37] CPU + V100 GPU
@@
[56] SHeGNN T, P, C, M, E Inference DGL T4 GPU
[52] SHeGNN T, P, M, E Inference PyG T4 GPU
@@ 4.1 Findings on Time Perspective
Time Complexity. T-1 The training time is related to the model’s time complexity. T-2 The time complexity of the message function used in the apply_edge substage affects the execution time of the aggregation stage [46, 48]. T-3 The time complexity for each vertex is affected by the dimensions of the input and output hidden features ($d_{in}$ and $d_{out}$) and the dimensions of the model parameters, such as the number of heads ($K$) in GAT [41] [46]. Accordingly, typical SHoGNNs are classified into four quadrants based on their time complexity of aggregation and combination, as shown in Fig. 4 [46]. T-4 All stochastic gradient descent based training algorithms suffer from exponential time complexity with respect to the number of layers [11, 48].
@@
Time Percentage. T-5 During SHoGNN training, the first layer consumes significantly more time than the second layer in model computation due to dimensionality reduction of vertex features and aggregation from more neighbors in the first layer [46]. T-6 During SHoGNN inference, both the gather and apply_vertex substages can dominate the execution time of each layer, varying across different datasets and models due to factors such as the variable dimensions of input and output vertex features [54]. T-7 Similarly, during SHoGNN training, the aggregation stage typically dominates the execution time of each layer in model computation, except in models with high computational complexity in the apply_vertex substage or graphs with low average vertex degrees [46]. T-8 Edge-related computations occupy most of the execution time of SHoGNN training and inference in most models [46]. For models with high edge computational complexity, most of time is spent on the messaging function for each edge. Otherwise, the aggregation of messages from all edges consume most of the time. T-9 There is no fixed hotspot in SHoGNN training, as substage time distribution varies greatly among models [61, 4].
T-10 Sampling in mini-batch training and inference of SHoGNNs constitutes a significant portion of total time [21, 27, 46, 18, 32, 2]. Sampling, typically performed on the CPU, can take up to 62% of an epoch’s time [21]. This bottleneck worsens if the CPU is attached to multiple GPUs and cannot produce enough samples to saturate them.
T-11 Mini-batch training for SHoGNNs faces time-consuming feature collection and data loading from CPU to GPU, particularly for loading vertex features [31, 29, 2]. The study [31] observed that 74% of time is spent on feature collection and data loading. This is mainly due to the large size of graph data, resulting in longer transfer time than model computation time. Moreover, using multiple GPUs within a single physical machine to speed up training proportionally increases the demand for data samples loaded from CPU to GPU. T-12 Furthermore, feature collection is CPU-intensive and takes significantly longer than CPU-GPU data movement [31]. When multiple GPUs are used, concurrent workers collecting features compete with samplers for CPU resources, resulting in an 88% increase in sampling time and a 59% increase in feature collection time compared to the single-GPU scenario [31]. T-13 Data loading time for training graph classification models constitutes a significant proportion of total training time [48]. Unlike the general data loading focus in T-4.1, T-4.1 specifically focuses on batching a set of small graphs into a single large batch.
T-14 The gradient synchronization in model computation becomes more time-consuming as the number of GPUs increases [29]. Unlike DNNs, GNNs have smaller model sizes with fewer layers and shared weights across all vertices. While the gradient synchronization is expected to occupy a small part of the execution time, it takes up a larger proportion as GPU count increases due to CPU-side imbalances, causing staggered GPU computations.
Relation between Time and Graph Characteristics. T-15 During SHoGNN inference, the execution time of the gather and apply_vertex substages in each layer increases linearly with the dimension of the output hidden features [53]. This is consistent with T-4.1. T-16 Additionally, the execution time of the apply_vertex substage is almost proportional to the dimension of the input vertex features [53]. This is consistent with T-4.1. T-17 Similarly to T-4.1, which focuses on SHoGNN inference, the execution time of the scatter, apply_edge, gather, and apply_vertex substages during SHoGNN training increases linearly with the dimension of the output hidden features [46]. This is consistent with T-4.1. T-18 The training and inference time of a layer is primarily affected by the dimensions of the input and output hidden features, increasing linearly with these dimensions [46]. This is consistent with T-4.1, T-4.1, T-4.1, and T-4.1. T-19 The time of the aggregation stage in a layer during SHoGNN training increases linearly with the average vertex degree of the graph [46]. T-20 The average vertex degree of graph affects the proportion of execution time for scatter, apply_edge, and gather substages in the total training time [46].
Relation between Time and Model Setup. T-21 Fixing other hyperparameters, each individual hyperparameter affects the training and inference time of a layer linearly [46]. T-22 Sampling reduces training time per batch only when the batch size is very small. As batch size increases, the overhead of sampling and data loading causes the training time to exceed that of full-batch training [46].
4.2 Findings on Parallelism Perspective
Parallelism Dependency. P-4 Inter-mini-batch parallelism in SHoGNN training is more complex than in DNN training due to the numerous dependencies between data samples [5]. P-5 Resource contention exists between different training stages when they run simultaneously [32]. Unlike DNN training, GNN training involves complex preprocessing tasks such as sampling and feature collection, which consume CPU, memory, and network bandwidth resources. Uncontrolled competition can cause poor performance, with some operations monopolizing resources and blocking others, preventing scaling with additional resources [32].
@@ 4.3 Findings on Compute Perspective
C-5 During SHoGNN training, computation is dominated by integer operations [4]. Results show that, on average, 64% of the executed instructions are integer instructions, while only 28.7% are floating-point instructions [4]. C-6 The scatter substage primarily involves data movement and address calculations for data accesses, preparing data for the subsequent substage [61, 40]. This explains the high proportion of integer operations observed in C-4.3.